Here is the question:
The units digit of a two-digit number is 2 less than the square of the tens digit. If 36 is added to the number, the result is the number with the digits reversed. Find the original number.
Thanks!

Answer by lwsshak3(11628)

The units digit of a two-digit number is 2 less than the square of the tens digit. If 36 is added to the number, the result is the number with the digits reversed. Find the original number.
**
let u=units digit
let t=tens digit
u=t^2-2
10t+u+36=10u+t
10t+(t^2-2)+36=10(t^2-2)+t
10t+t^2-2+36=10t^2-20+t
10t+t^2+34=10t^2-20+t
9t^2-9t-54=0
t^2-t-6=0
(t-3)(t+2)=0
t=-2 (reject)
or
t=3
u=t^2-2=9-2=7
original number: 37
Check: 37+36=73